Consider the matrices A =   3 0 2 4 −6 3 −2 1 8   ,B =   −5 1 1 0 3 0 7 6 2   , C =   1 1 1 2 3 −1 3 −5 −7   , Verify the following expressions (where possible and give reasons) (i) A + (B + C) = (A + B) + C and A(BC) = (AB)C. (ii) (a − b)C = aC − bC and a(B − C) = aB − aC, where a = −2, b = 3 . (iii) (A T ) T = A and (A − B) T = A T − B T .


1
Expert's answer
2021-05-10T13:11:10-0400

Given matrices-


"A=\begin{bmatrix}
 3 & 0&2 \\
 4 & -6&3\\-2&1&8
\end{bmatrix},B=\begin{bmatrix}
 -5 & 1&1 \\
 0 & 3&0\\7&6&2
\end{bmatrix},C=\begin{bmatrix}
 1 & 1&1 \\
 2 & 3&-1\\3&-5&-7
\end{bmatrix}"


"(i) LHS- A+(B+C)"


"=\begin{bmatrix}
 3 & 0&2 \\
 4 & -6&3\\-2&1&8
\end{bmatrix}+(\begin{bmatrix}
 -5 & 1&1 \\
 0 & 3&0\\7&6&2
\end{bmatrix}+\begin{bmatrix}
 1 & 1&1 \\
 2 & 3&-1\\3&-5&-7
\end{bmatrix})\\[9pt]=\begin{bmatrix}
 3 & 0&2 \\
 4 & -6&3\\-2&1&8
\end{bmatrix}+\begin{bmatrix}
 -4& 2&2 \\
 2 & 6&-1\\10&1&-5
\end{bmatrix}=\begin{bmatrix}
 -1 & 2&4 \\
 6 & 0&2\\8&2&3
\end{bmatrix}"


RHS-(A+B)+C

"=(\begin{bmatrix}
 3 & 0&2 \\
 4 & -6&3\\-2&1&8
\end{bmatrix}+\begin{bmatrix}
 -5 & 1&1 \\
 0 & 3&0\\7&6&2
\end{bmatrix})+\begin{bmatrix}
 1 & 1&1 \\
 2 & 3&-1\\3&-5&-7
\end{bmatrix}\\[9pt]=\begin{bmatrix}
 -2 & 2&2 \\
 4 & -3&3\\5&7&10
\end{bmatrix}+\begin{bmatrix}
 1 & 1&1 \\
 2 & 3&-1\\3&-5&-7
\end{bmatrix}=\begin{bmatrix}
 -1 & 2&4 \\
 6 & 0&2\\8&2&3
\end{bmatrix}"


Hence LHS=RHS.


"(ii) Lhs= (a-b)C=(-2-3)C=-5C=-5\begin{bmatrix}
 1 & 1&1 \\
 2 & 3&-1\\3&-5&-7
\end{bmatrix}" ="\begin{bmatrix}
 -5 & -5&-5 \\
 -10 & -15&5\\-15&25&49
\end{bmatrix}"


"RHS=aC-bC=-2C-3C=-2\begin{bmatrix}
 1 & 1&1 \\
 2 & 3&-1\\3&-5&-7
\end{bmatrix}-3\begin{bmatrix}
 1 & 1&1 \\
 2 & 3&-1\\3&-5&-7
\end{bmatrix}\\[9pt]=\begin{bmatrix}
 -2 & -2&-2 \\
 -4 & -6&2\\-6&10&14
\end{bmatrix}-\begin{bmatrix}
 3 & 3&3 \\
 6 & 9&-3\\9&-15&-21
\end{bmatrix}\\[9pt]=\begin{bmatrix}
 -5 & -5&-5 \\
 -10 & -15&5\\-15&25&49
\end{bmatrix}"


Hence, RHS=LHS


"(iii) (A^T)^T=\begin{bmatrix}
 3 & 0&2 \\
 4 & -6&3\\-2&1&8
\end{bmatrix}''=\begin{bmatrix}
 3 & 4&-2 \\
 0& -6&1\\2&3&8
\end{bmatrix}'=\begin{bmatrix}
 3 & 0&2 \\
 4 & -6&3\\-2&1&8
\end{bmatrix}=A"


Hence, "(A^T)^T=A"